Operations Report – April 24th

Crew 211 Operations Report 24-APR-2019

SOL: 3

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10:30

On (last night): 8:30 pm

Off (this morning): 7 am

Solar— SOC 100 % at 20:10 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons

Water (auxillary tank) – 0 gallons

Water (static tank) – ~240 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 110 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 6 units

Toilet tank emptied: yes

Deimos rover used: no (rover is out of order according to Director)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: yes

Hours: 89.0h

Beginning charge: 100

Ending charge: 84.4

Currently charging: yes

Opportunity rover used: yes

Hours: 57.6

Beginning charge: 95

Ending charge: 74

Currently charging: no

Curiosity rover used: yes

Hours: 90.8h

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Notes on rovers: Yesterday during the EVA the rover Opportunity suffered overheating and could not be started back. This morning’s mission was to bring it back so that the mechanics could fix it.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Some basil and tomatoes were harvested for cooking.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Simon finished his UWB stands and Carl continued assembling more probes during the afternoon after children finished visiting us.

Summary of any observatory issues: The Astronomer’s computer encounters sometimes freeze and displays artifacts on the screen. This might be the beginning of a GPU failure.

Summary of health and safety issues: The member with the twisted ankle feels better and can walk, but for extra safety HSO does not recommend EVA.

Questions, concerns and requests to Mission Support:

* We had an Internet problem last night which prevented us to send the operations report to CapComm. It solved itself at around 10am this morning.

* We overestimated the static water tank capacity, and after checking the level properly with a rope and measuring volumes we reestimated its current level. We will try to diminish consumption as much as we can, but it is already very low and we are investigating other reasons.

* The water pump leak is at a plastic joint. To fix it we would need to stop the waterflow and replace the joint with a new one. We will try to document further the following days. Attaching a picture of the pump with the mail (faulty joint is in red, and you can see that water has been leaking for it).

EVA Report – April 24th

Crew 211 EVA Report 24-04-2019

EVA #3

Author: Simon Collignon

Purpose of EVA: Taking back the rover opportunity to the MDRS and exploring the surroundings of the hab.
Start time:09:00 am
End time:10:30 am

Narrative: The aim of this EVA was to reach the rover opportunity which had a battery failure yesterday on Cow Dung Road in order to drive it back home. After the rover’s rescue, we’ve explored the surroundings of the MDRS in order to have a better understanding of the geology near the Hab.

Destination: Intersection of Stream Bed Connector and Cow Dung Road, Hab View North.

Coordinates (use UTM NAD27 CONUS): Latitude: 4251500, longitude: 519000 (Intersection of Stream Bed Connector and Cow Dung Road)
Latitude: 4250900, longitude: 517900 (Hab View North)

Participants: Carl-Henrik Dahlqvist (commander of the EVA), Simon Collignon (medical officer), Chloe Peduzzi.
Road(s) and routes per MDRS Map: On Cow Dung Road from Hab to Stream Bed Connector then on Cow Dung Road from Stream Bed Connector to Hab. After parking the rovers at the hab, we walked around the MDRS towards North and directly turned towards West in order to reach the Hab View North.

Mode of travel: Driving and hiking

Greenhab Report – April 24th

Crew 211 GreenHab Report – 24-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open + fan turned up to 3 between 10:00 and 18:00 + fan turned up to 2 the rest of the time

Shade cloth: 80%

Average temperatures:

Low: 14°C

High: 44°C

Hours of supplemental light: 0

Daily water usage for crops: 3 gallons

Water in Blue Tank: 105 gallons

Time(s) of watering for crops: 14:00

Changes to crops: All plants still look healthy. More flowers are visible on the tomato plant of the GreenHab. Lettuces (old seed) produce flowers. Romaine lettuce still grow very well. We have harvested the 2 plants of Lettuce Red Leaf as they wouldn’t grow anymore.

Narrative: It was hotter today despite the fan was turned up to 3. Nevertheless I will do the same tomorrow. For our dinner we have harvested lettuce, parsley and basil (from the GreenHab) and tomatoes (from the plants we brought). I have sow bean and tomato seeds according to their respective treatments for my experiment.

Harvest: 11.4g basil + 21,2g lettuce + 5,6g parsley (+ 50g tomatoes from plants brought by our crew)

Support/supplies needed: None

Astronomy Report – April 24th

Astronomy Report

Name: Lieffrig Eléonore

Crew: 211
Date: 24 April 2019


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: 3C273

Images submitted with this report: None

Problems Encountered: None


Solar Features Observed:

Granules, prominences at the surface of the sun.
We were really happy to be able to see the eruptions on the pictures you’ll find enclosed.

Problems Encountered: The tracking of the sun doesn’t work, I have to follow the sun manually, could you help please ?

Science Report – April 24th

Science Report 24 April 2019 Crew 211 UCL to Mars Submitted by Science Officer Maxime Bernard 1) Commander Carl-Henrik Dahlqvist The 8 CubeSat cases are assembled and two of them are already functional. He is working on the code to increase the size of the cubelanders network. 2) XO Officer All four antennas are assembled and fixed to the poles. At the moment, he is still working on the software side of the project. 3) HSO Officer Benjamin Flasse Every morning and every evening, the HSO still gather data about the mensurations, weight, brain awakeness, tension and pulse of the crew. In the meantime he is also making an inventory of all the first aid kits to ensure they are all complete and reliable. 4) Crew Astronomer Eleonore Lieffrig Our astronomer managed to take pictures of the sun with the telescope and solar eruptions could be seen. She is also preparing the observation of a quasar using the remote controlled telescope. 5) Crew Engineer Julien Amalaberque No major milestone reach for now. 6) Crew Biologist Chloé Peduzzi The crew engineer and the XO managed to inverse the transformer to allow Chloé to use her LED light for her first sample of spirulina. She also placed a temperature sensor near the spirulina to make sure the temperature will never exceed 30°C. She then started her second experiment in which she is going to plant tomato and bean seeds in different medium composed of different quantities of compost/martian soil with mycorrhiza and/or hydrogel which are going to be dry or not. It will finally constitute a total of 16 different medium. 7) Crew Journalist Nathan Pechon We made a meeting during which he presented the voting rules and the different possible regimes available according to the talks we all had with him earlier. During this meeting we voted for the political regime we thought would suit the best an early martian colony. 9) Crew Scientist Maxime Bernard The detector finally arrived in Hanksville and Maxime together with Shannon picked it up around 6:30 pm.


Crew 211 Journalist Report 24Apr2019

[title Journalist Report – April 24th]

Today was our third simulation day. We woke up at 7:30 am and we had breakfast at 8:00. Nathan and Chloé prepared choco Panecakes. At 9 am, a part of the crew had to go in EVA to rescute Opportunity rover. After that, they walked around the space station during one hour.

At the same time, Benjamin did a medical inventory. Eleonore did pictures of the sun with the solar telescope.

At 10:30 am, students from Colorado visited the MDRS. The children were very interested by the station. We showed all rooms of the Hab, the greenhab, the Ram, science dome and the solar telescope.

Julien prepared the lunch. We ate rice with rest of the apple cake. We ate apple cake for dessert. After meal, each of us worked on his experiences.

At 6 pm, we prepared the Cap Com. Each of us is working on his report and is going to send it between 7 to 9 pm. After the Cap Com, we will eat pita with salad and tomatoes from the greenhab. Mmmmmmmmm……

Like usual, we will do a crew meeting to speak about the day.

Sol Summary – April 24th

Crew 211 Sol 3 Summary Report 24-APR-2019

Sol: 3

Summary Title: A day with Martians

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: After the daily medical check-up and a nice breakfast with chocolate pancakes, members of the crew left at 9:00 am for an EVA to get back Opportunity still stuck near Candor Chasma. To avoid any other issues with the rover, the EVA team got directly back to the Hab and continued its EVA on foot to explore the hills surrounding the base. The surprise of today came at 10:15 am when we learned that the student visit scheduled for tomorrow would in fact start 15 minutes later. We spent the next four hours with two groups of students visiting the different infrastructures of the MDRS and presenting our research works. These four hours have been very enriching for both the students and the UCL to Mars team. After the lunch, we all worked on our experiments. Tonight will be dedicated to sky observations with the remote controlled telescope.

Look Ahead Plan: An EVA is planned in the morning to explore Special Region. Half of the crew will stay at the base to work, as usual, on their experiments. The rest of the day will be mainly devoted to these experiments.

Anomalies in work: No issues.

Weather: Hot, bright and sunny all day.

Crew Physical Status: Twisted ankle still bloated (See HSO reports)

EVA: The aim of this third EVA was to recover the Opportunity rover. Once recovered, we decided to explore on foot the hills surrounding the MDRS.

Reports to be filed: Journalist Report, GreenHab Report , HSO report, Science Report, Operations Report, Astronomy Report, EVA report, EVA Request, Daily Photo Report + Timelapse

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site